# Using Biot-Savart to shrink Eulerian domains while maintaining or improving external flow accuracy

### v1

@@ 6 Conclusion
In this manuscript, we presented a novel boundary condition formulation for the incompressible Navier-Stokes equations. The methods rely on a Biot-Savart integral of the vorticity inside the domain to set the velocity boundary conditions on the domain’s boundaries. We couple these boundary conditions with an explicit projection scheme to solve the momentum and continuity equation in their primitive variables form. Leveraging the fast decaying properties of the Biot-Savart kernel, we use a multilevel algorithm to reduce the number of operations required to update the boundary nodes from $O(N)$ to $O(\log N)$ and show that the error introduced by this multilevel approach is bounded and can be estimated. We demonstrate that when a body is immersed in a domain using these boundary conditions, a coupling between the immersed body and the domain’s boundary is introduced. This coupling only results in a weakly non-linear pressure Poisson equation and a simple differed correction approach can be used to solve the coupled projection/Biot-Savart step, allowing the use of standard Poisson solver.
@@
Surprisingly, for vortex street applications, the method demonstrates its robustness and perfectly captures the drag forces and vortex shedding frequencies generated by a 2D circular cylinder at $Re=200$, even with minimal domain sizes ($5D\times 3D$). Classical reflection boundary conditions requires a domain at least 10 times bigger to generate time-varying forces indistinguishable from that of the domain with our Biot-Savart boundary conditions. Lastly, we show that the method is limited when strong wake history influences the flow in the near wake. For a high-amplitude, high-Strouhal number heaving airfoil, we show that capturing the deflected wake of the flow requires significantly larger domain (with our Biot-Savart boundary conditions) and that the computational advantage is somewhat lost.
These results show that our novel boundary conditions are very effective for confined vorticity applications, such as accelerated bodies in the initial phase ($t^{*}<10$), when the vorticity is still confined to the body, or flow with very weak wake dependency.

### v2

@@ 6 Conclusion
In this manuscript, we presented a novel boundary condition formulation for external incompressible flow simulations. The method relies on a Biot-Savart integral of the vorticity inside the domain to set the velocity boundary conditions on the domain boundaries. We couple these boundary conditions with an explicit projection scheme to solve the momentum and continuity equation in their primitive variables form. Leveraging the fast decaying properties of the Biot-Savart kernel, we use a multilevel algorithm to reduce the number of operations required to update the boundary nodes from $O(N)$ to $O(\log N)$ and show that the error introduced by this multilevel approach is bounded and can be estimated. We demonstrate that when a body is immersed in a domain using these boundary conditions, a coupling between the immersed body and the domain’s boundary is introduced. This coupling only results in a weakly non-linear pressure Poisson equation and a simple differed correction approach can be used to solve the coupled projection/Biot-Savart step, allowing the use of standard Poisson solver.
@@
Surprisingly, for vortex street applications, the method demonstrates its robustness and perfectly captures the drag forces and vortex shedding frequencies generated by a 2D circular cylinder at $\text{Re}=200$, even with minimal domain sizes ($5D\times 3D$). Classical reflection boundary conditions require a domain 48 times bigger to generate time-varying forces indistinguishable from that of the domain with our Biot-Savart boundary conditions. Lastly, we show that the method is limited when strong wake history influences the flow in the near wake. For a high-amplitude, high-Strouhal number heaving airfoil, we show that capturing the deflected wake of the flow requires significantly larger domain, implying that the computational advantage of the new boundary condition is reduced.
The generality and accuracy of the new approach, even when simulating flows with vortex wakes that are truncated by the domain boundaries, opens a huge range of potential applications. In addition to the welcomed speed-up enabled by using frankly tiny computational domains, the continued accuracy of the simulations challenges the notion of the importance of far-field wake dynamics on the near-body flow, providing compelling evidence to the contrary.
Appendix A Accounting for symmetries in the Biot-Savart integral
We briefly describe how symmetries can be accounted for in the multilevel evaluation of the Biot-Savart integral Eq. 8. We focus on a description of the 2D case; the extension to the 3D is trivial. For a flow with symmetries, like the 2D flow around a flat plate, see Figure 13, whose axis of symmetries is represented by the dash-dotted line, the computational domain $\Omega$ only models the upper part of the plate. The domain $\Omega^{\dagger}$ is the image of the computational domain $\Omega$ about the symmetry axis.
Evaluation of the induced velocity at a point on the boundary of the computational domain $\vec{x}\in\partial\Omega$ (apart from points on the symmetry plane which have the reflection condition applied to them) on the $i$th-level is given by
${\vec{u}}({\vec{x}})=\int_{\mathcal{R}^{(i)}}K_{n}({\vec{x}}-\vec{y})\times\hat{e}_{z}\omega({\vec{y}})\text{ d}\vec{y}+\int_{{\mathcal{R}^{(i)}}^{\dagger}}K_{n}({\vec{x}}-\vec{y}^{\dagger})\times\hat{e}_{z}\omega^{\dagger}({\vec{y}}^{\dagger})\text{ d}\vec{y}^{\dagger}$ (28)
where $\mathcal{R}^{(i)}\equiv\mathcal{D}^{(i)}\cap\Omega$ and ${\mathcal{R}^{(i)}}^{\dagger}\equiv\mathcal{D}^{(i)}\cap\Omega^{\dagger}$. Additionally, $\overline{\mathcal{R}^{(i)}}^{\dagger}\subset\mathcal{R}^{(i)}$ is the images of ${\mathcal{R}^{(i)}}^{\dagger}$ in $\Omega$ and ${\vec{y}}^{\dagger}$ is the images of $\vec{y}$ in $\Omega^{\dagger}$, respectively.
Points inside ${\mathcal{R}^{(i)}}$ (blue region) are evaluated on the computational domain $\Omega$. For points in ${\mathcal{R}^{(i)}}^{\dagger}$ (orange region), the image vorticity is computed by applying the symmetric condition to the vorticity in $\overline{\mathcal{R}^{(i)}}^{\dagger}$ (pink region)
$\omega^{\dagger}(\vec{y}^{\dagger})=-\omega(\vec{y}),\quad\quad\vec{y}\in\overline{\mathcal{R}^{(i)}}^{\dagger},$ (29)
for 2D flows; for 3D flows, computing the image vorticity vector requires computing its reflection about the symmetry plane. Substituting in Eq. 28 yields an integral only over the region $\mathcal{R}^{(i)}$
${\vec{u}}({\vec{x}})=\int_{\mathcal{R}^{(i)}}\left(K_{n}({\vec{x}}-\vec{y})-K_{n}({\vec{x}}-\vec{y}^{\dagger})\right)\times\hat{e}_{z}\omega({\vec{y}})\text{ d}\vec{y},$ (30)
where the second term in the parenthesis only appears if $\mathcal{D}^{(i)}$ overlaps with the domain’s image $\Omega^{\dagger}$. A similar expression is found for 3D flows. As Eq. 30  solely relies on the vorticity in the computational domain $\Omega$, no additional computations are required, and accounting for the image’s induced velocity simply requires accounting for its influence through the kernel $K_{n}$.
$\mathcal{D}^{(i)}$
$\Omega$
$\vec{y}$
$\vec{y}\,^{\dagger}$
$\vec{x}$
${S}\,^{(i)}$
$\mathcal{R}^{(i)}$
${\mathcal{R}^{(i)}}^{\dagger}$
$\overline{\mathcal{R}^{(i)}}^{\dagger}$
${S}\,^{(i)}$
$\Omega^{\dagger}$
$\partial\Omega$
Figure 13: Schematic of the approach used to account for symmetries in the multilevel evaluation of the Biot-Savart integral. The induced velocity at a point $\vec{x}$ includes the contribution from $\vec{y}$ and its image point $\vec{y}^{\dagger}$. The computational domain $\Omega$ and its image $\Omega^{\dagger}$ intersect with a multilevel domain $\mathcal{D}^{(i)}$ of half-width $S^{(i)}$ producing the regions $\mathcal{R}^{(i)}$, ${\mathcal{R}^{(i)}}^{\dagger}$ and $\overline{\mathcal{R}^{(i)}}^{\dagger}$.